OpenStudy (anonymous):

could someone help me quick Which of the following values "completes the square," or creates a perfect square trinomial, for x2 + 10x + ___? 25 10 5 –5

OpenStudy (michele_laino):

hint: what is \((x+5)^2=...?\)

OpenStudy (anonymous):

add square of half the coefficient of x

OpenStudy (anonymous):

thank you @Michele_Laino could you help with another one by chance

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

Rewrite f(x) = –2(x − 3)2 + 2 from vertex form to standard form. f(x) = –2x2 − 18 f(x) = –2x2 + 12x + 20 f(x) = –2x2 + 12x − 16 f(x) = 4x2 − 24x + 38

OpenStudy (anonymous):

@Michele_Laino ^

OpenStudy (michele_laino):

here we have to compute the square of binomial, so we can write this: \[f\left( x \right) = - 2\left( {{x^2} + 9 - 6x} \right) + 2 = ...?\] please continue

OpenStudy (anonymous):

so i got f(x)= -2x^2 -18 +12x +2 ? is that right

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

yes! Now you have to simplify again, so what is \(-18+2=...?\)

OpenStudy (anonymous):

-16 so it would be answer choice b? @Michele_Laino

OpenStudy (michele_laino):

you have to search for the option which contains \(-16\)

OpenStudy (anonymous):

thats what i did. sorry to ask but do you have time for one more by chance @Michele_Laino

OpenStudy (michele_laino):

I think it is option C. Yes! I can help you!

OpenStudy (anonymous):

@Michele_Laino that what i meant i typed the wrong letter Solve x2 + 8x − 3 = 0 using the completing-the-square method.

OpenStudy (michele_laino):

If I add and subtract 16, I get this: \[\Large \begin{gathered} {x^2} + 8x - 3 = {x^2} + 8x + 16 - 16 - 3 = \hfill \\ \hfill \\ = \left( {{x^2} + 8x + 16} \right) - 16 - 3 = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

I am very confused on this where did you get the 16 from @Michele_Laino

OpenStudy (michele_laino):

since the three terms inside the parentheses are the square of a binomial. What is such binomial?

OpenStudy (anonymous):

@Michele_Laino i am sorry but i still am not getting this

OpenStudy (michele_laino):

please compute the square of this binomial: \((x+4)^2=...?\)

OpenStudy (anonymous):

i got x^2 + 8x + 16 thats where you got the 16 right? @Michele_Laino

OpenStudy (michele_laino):

yes! That is why I chose \(16\)

OpenStudy (anonymous):

ok so so how do i solve this though @Michele_Laino

OpenStudy (michele_laino):

so we can rewrite your original equation as below: \[\begin{gathered} {x^2} + 8x - 3 = {x^2} + 8x + 16 - 16 - 3 = \hfill \\ \hfill \\ = \left( {{x^2} + 8x + 16} \right) - 16 - 3 = \hfill \\ \hfill \\ = {\left( {x + 4} \right)^2} - 19 \hfill \\ \end{gathered} \] therefore: \[{x^2} + 8x - 3 = 0 \Rightarrow {\left( {x + 4} \right)^2} - 19 = 0\]

OpenStudy (michele_laino):

In other words the original equation, is equivalent to this one: \[{\left( {x + 4} \right)^2} - 19 = 0\]

OpenStudy (anonymous):

my answer choices have an addition sign over a subtraction sign so what would the answer be @Michele_Laino

OpenStudy (michele_laino):

If we add \(19\) to both sides, we get: \[{\left( {x + 4} \right)^2} - 19 + 19 = 0 + 19\] please combine similar terms

OpenStudy (anonymous):

ok now i understand i know i sound needy but is there anytime for 1 more i just dont understand it @Michele_Laino

OpenStudy (michele_laino):

ok! I can help you :)

OpenStudy (anonymous):

Using the completing-the-square method, rewrite f(x) = x2 − 8x + 3 in vertex form. f(x) = (x − 8)2 f(x) = (x − 4)2 − 13 f(x) = (x − 4)2 + 3 f(x) = (x − 4)2 + 16

OpenStudy (anonymous):

@MissViolet

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

again I add and subtract \(16\) at the right side, so I get this: \[\begin{gathered} f\left( x \right) = {x^2} - 8x + 3 = {x^2} - 8x + 16 - 16 + 3 = \hfill \\ \hfill \\ = \left( {{x^2} - 8x + 16} \right) - 16 + 3 = \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

the three terms inside the parentheses are the square of a binomial. Do you know what is such binomial?

OpenStudy (anonymous):

no i dont @Michele_Laino

OpenStudy (michele_laino):

please, compute this: \[{\left( {x - 4} \right)^2} = ...?\]

OpenStudy (anonymous):

x^2 - 8x + 16 ?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

correct! so we can rewrite your function as below: \[\begin{gathered} f\left( x \right) = {x^2} - 8x + 3 = {x^2} - 8x + 16 - 16 + 3 = \hfill \\ \hfill \\ = \left( {{x^2} - 8x + 16} \right) - 16 + 3 = \hfill \\ \hfill \\ = {\left( {x - 4} \right)^2} - 16 + 3 \hfill \\ \end{gathered} \] so, what is the right option?

OpenStudy (anonymous):

Answer choice b @Michele_Laino

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

thank you so much @Michele_Laino next time i need help ill come to you again

OpenStudy (michele_laino):

ok! :)

OpenStudy (anonymous):

btw for anyone going to use this they are all correct